40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12957     Entry time: Fri Apr 28 19:32:06 2017     In reply to: 12954
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking - PRCL angular FF 

I took a closer look at the POP QPD/ PRC angular feedforward situation yesterday. I thought it would be useful to have a POP QPD MEDM screen. Looking at the PIT and YAW channel filter modules, the anti-whitening filters seemed different from what we have for other channels that are connected to the Pentek interface board (e.g. MCL). So I copied over the 150:15 (z:p) filter, and also turned on a 60Hz comb. The LSC offsets script does not set the dark offsets for this QPD, so I manually put in the dark offsets for the PIT, YAW and SUM channels as well. For the locking, I first locked the arms on IR an dither aligned them. Then I locked the PRMI on carrier, ran the PRC dither alignment, and went over to the ITMX pickoff table and centered the beam on the QPD by making the PIT and YAW channel timeseries oscillate around approximately zero. 

After these tweaks, I collected ~40mins of data with the angular FF OFF/ON. I did not DC couple the ITM Oplev servos, but Eric tells me that this did not make a difference to the achievable subtraction in the past. Here is the frequency domain multicoherence analysis - I used the BS_X and BS_Y seismometer channels as witnesses. I've also put a plot with what the raw FF filter coefficients look like (no fitting yet). 

      

Looks like we can do better for both DOFs - it even seems like we are injecting noise with the current FF filters in some bands, perhaps we can do a better job of rolling off the filters outside the band of interest. Eric and I were discussing MATLAB's "reduce" routine for this purpose, I will play around with it and see if I get a better fit.

Unfortunately, I encountered a strange error when trying to pull data with nds2 today, it kept complaining RuntimeError: Too many channels or too much data requestedeven though I have pulled longer stretches of data for more channels with 16k sampling rate as recently as last week. Shorter duration requests (<600 seconds) seemed to work fine though... So I had to use cds.getdata to pull the data, and they're much too large to attach. Has anyone else encountered a similar error?


The mystery of the spots on the ITMs when the PRC is locked on carrier remains - after talking this over with Koji, we figured that even with the carrier resonant, the spot will be much dimmer than the spots when the arms are locked, but what I see on the cameras is still a pretty beefy spot. The real cavity mode is actually visible where it should be (I marked the locations of the spots with arms well-aligned with a marker on the monitors), as given away by some twinkling that is visible only when the cavity is locked. But what ghost beam is so intense it looks almost as bright as when the arm is locked?

GV 10pm 28 April 2017: Turns out this is the spot from the single bounce off the ETM transmitting back through the ITM and hitting the suspension cage (hence the bright spot). Johannes and I confirmed by moving the ETM, the spot moved with it. I just never paid attention to this spot before.

Attachment 1: PRC_angularFF.pdf  389 kB  | Hide | Hide all
PRC_angularFF.pdf
Attachment 2: PRC_TFs.pdf  391 kB  | Hide | Hide all
PRC_TFs.pdf
ELOG V3.1.3-